# Simultaneous global and local clustering in multiplex networks with covariate information

### v1

@@ 1 Introduction
Modern statistics has seen a growing interest in the use of network models to mathematically describe complex relationships between a large number of connected entities. Network data is often highly complex, with multiple types of connections existing between the nodes in the network. To model such heterogeneity of connection types, multiplex networks have been developed as compact representations, with multiple layers being considered simultaneously, each describing a different connection type between its nodes. Examples of applications include neuroimaging, where regions of interest in the brain represent nodes and layers correspond to different patients (Paul and Chen, 2018); actions of individuals on a social media platform, where layers might represent “follow” or “retweet” interactions (Greene and Cunningham, 2013); or economic data, where layers capture different types of financial exposure (Poledna et al., 2015). For a survey of multiplex network applications, see, for example, Kivelä et al. (2014).
@@
There has been work to extend single-layer clustering methodology to the multiplex setting, much of which considers variants of a multiplex SBM. Broadly speaking, the approaches can be divided into those that attempt primarily to cluster the layers, and potentially the nodes therein (Stanley et al., 2015; Reyes and Rodriguez, 2016; Paul and Chen, 2018; Jing et al., 2021; Young et al., 2022; Josephs et al., 2023; Zheng et al., 2024), and those that seek only to infer either an across-layer community structure, or within-layer node groups without clustering the layers (Amini et al., 2024; Berlingerio et al., 2011; Kuncheva and Montana, 2015). To the best of our knowledge, a methodology that provides both an across-layer, or global, clustering of each node alongside a layer-level, or local, clustering has not yet been proposed in the literature. In this paper, we aim to address this gap by developing a framework for hierarchical clustering of nodes in a multiplex network that allows heterogeneous clustering across layers, infers latent groups at local and global levels, and incorporates nodal covariate information to guide clustering. This particular type of hierarchical clustering finds relevant applications across several fields. In economics, global clustering can describe trade patterns between entities (Fagiolo et al., 2010). In biology, it helps understand how genes or proteins participate in different biological pathways while also inferring overarching functional groupings (Barabasi and Oltvai, 2004).
@@
With the exception of Reyes and Rodriguez (2016); Josephs et al. (2023) and Amini et al. (2024), most existing methods assume a known number of communities for layers and nodes. Furthermore, to the best of our knowledge, there are no existing methodologies that have the flexibility of simultaneously allowing for both an unknown number of communities and nodal covariate information. Approaches for dealing with an unknown number of nodes usually rely on Bayesian nonparametric priors; for example, Reyes and Rodriguez (2016) modify the infinite relational model of Kemp et al. (2006) to cluster networks with similar community structure. Similarly, the hierarchical stochastic blockmodel (HSBM; Amini et al., 2024) employs a hierarchical Dirichlet process (HDP; Teh et al., 2006) to infer layer-level communities whilst sharing information between the layers. The method proposed in this work is closely related to the nested stochastic blockmodel (NSBM) of Josephs et al. (2023), which uses a nested Dirichlet process (NDP; Rodríguez et al., 2008) to provide a clustering of layers and of the nodes within those layers. The nonparametric prior allows for the number of layer communities and communities within layers to be unspecified and automatically selected. However, the NSBM does not provide a global clustering of nodes, nor is it adapted for the inclusion of nodal covariate information, which we address within the model proposed in this work.
@@ 2 Hierarchical multiplex stochastic blockmodels
We consider a directed multiplex network with $L$ layers and $N$ nodes. We assume that all layers share the same node set $\mathcal{V}:=[N]$, where we write $[m]=\{1,\dots,m\},\ m\in\mathbb{N}$. Each layer $\ell\in[L]$ consists of a graph $\mathcal{G}_{\ell}=(\mathcal{V},\mathcal{E}_{\ell})$, where $\mathcal{E}_{\ell}\subseteq\mathcal{V}\times\mathcal{V}$ is a layer-specific edge set, and we say that $(i,j)\in\mathcal{E}_{\ell}$, $i,j\in\mathcal{V}$, if nodes $i$ and $j$ form a connection of type $\ell$. The full multiplex network is denoted by $\mathcal{G}=(\mathcal{V},\{\mathcal{E}_{1},\dots,\mathcal{E}_{L}\})$. Each graph $\mathcal{G}_{\ell}$ could equivalently be represented by its adjacency matrix $A_{\ell}=\{A_{\ell ij}\}_{i,j=1}^{N}\in\{0,1\}^{N\times N}$, where $A_{\ell ij}=\mathbb{I}_{\mathcal{E}_{\ell}}\{(i,j)\}$ and where $\mathbb{I}_{\cdot}\{\cdot\}$ denotes the indicator function. The multiplex network $\mathcal{G}$ can then be represented by an adjacency tensor $A=\{A_{\ell ij},\ \ell\in\{1,\dots,L\},\ i,j\in\{1,\dots,N\}\}\in\{0,1\}^{L\times N\times N}$.
We model the behaviour of each layer by conditionally independent stochastic blockmodels (SBMs; Holland et al., 1983). To parameterise the SBM of layer $\ell\in[L]$, we assume that each node $i\in\mathcal{V}$ belongs to a group $z_{\ell i}\in\mathbb{N}$. Note that, in a Bayesian nonparametric fashion, we consider the number of clusters to be not fixed a priori, and it is instead treated as a random quantity that can grow with the number of nodes. Given these layer-level group $z_{\ell i},\ \ell\in[L],\ i\in[N]$, the adjacency matrix $A_{\ell}$ is modelled as
@@
and we write $\gamma=\{\gamma_{ks}\}$ and $\gamma^{\prime}=\{\gamma^{\prime}_{ks}\}$ for the matrices of storing parameters. Additionally, we assume that a vector $\bm{x}_{i}\in\mathbb{R}^{P}$ of covariates, which is assumed to be completely observed, is associated to each node $i\in\mathcal{V}$. We use the covariates $\bm{x}_{i}$ to construct a the node-specific probability vectors $\bm{\tau}_{i}$ in (2.3) for each $i\in\mathcal{V}$. To allow for learning of an unbounded number of communities, we utilise a stick-breaking representation (Sethuraman, 1994) for each $\bm{\tau}_{i}$, which ensures that $\sum_{k=1}^{\infty}\tau_{ik}=1$, where $\tau_{ik}$ is the probability that node $i$ belongs to global group $k$. In particular, we employ the probit stick-breaking construction of Chung and Dunson (2009) and Rodriguez and Dunson (2011). In this model, $\bm{\tau}_{i}=(\tau_{i1},\tau_{i2},\dots)^{\intercal}$ is obtained as
@@
$A_{\ell ij}\mid z_{\ell i},z_{\ell j},\rho_{z_{\ell i}z_{\ell j}}$ $\sim\text{Bernoulli}(\rho_{z_{\ell i}z_{\ell j}}),$ for $\text{for all }i,j\in\mathcal{V},\;i\neq j,\;\ell\in[L],$ (2.6)
$\rho_{km}$ $\sim\text{Beta}(\alpha_{0},\beta_{0}),$ for all $\text{for all }k,m\in\mathbb{N},$ (2.7)
$z_{\ell i}\mid w_{i},\bm{\gamma}_{w_{i}}$ $\sim\text{Categorical}(\bm{\gamma}_{w_{i}}),$ for $\text{for all }i\in\mathcal{V},\;\ell\in[L],$ (2.8)
$w_{i}\mid\bm{\tau}_{i}$ $\sim\text{Categorical}(\bm{\tau}_{i}),$ for all $\text{for all }i\in\mathcal{V},$ (2.9)
$\tau_{ik}$ $=\Phi\{\bm{x}_{i}^{\intercal}\bm{\varphi}_{k}\}\prod_{\ell=1}^{k-1}\left(1-\Phi\left\{\bm{x}_{i}^{\intercal}\bm{\varphi}_{\ell}\right\}\right),$ for $\text{for all }i\in\mathcal{V},\ k\in\mathbb{N},$ (2.10)
$\bm{\varphi}_{k}\mid\bm{\varphi}^{0}_{k},\sigma_{k}^{2}$ $\sim\text{Normal}(\bm{\varphi}^{0}_{k},\sigma_{k}^{2}I_{P}),$ for all $\text{for all }k\in\mathbb{N},$ (2.11)
$\bm{\varphi}^{0}_{k}$ $\sim\text{Normal}(\bm{\mu},I_{P}),$ for all $\text{for all }k\in\mathbb{N},$ (2.12)
$\sigma_{k}^{2}$ $\sim\text{Inverse\text{-}Gamma}(\nu_{0},\omega_{0}),$ for all $\text{for all }k\in\mathbb{N},$ (2.13)
$\bm{\gamma}_{k}$ $\sim\text{GEM}(\eta_{0}),$ for all $\text{for all }k\in\mathbb{N}.$ (2.14)
@@
Figure 2: Toy example of the group structure of nodes in a multiplex network with $L=3$ layers. There are two global groups, indicated by the shape of the nodes, and three layer groups, indicated by the colours. Probabilities of layer group assignments given the global groups are indicated in the row vectors.
@@
On the other hand, the NSBM is used in Josephs et al. (2023) for clustering of layers in a multiplex network where each layer follows an SBM. Layers are assigned to a community which determines its connection probability matrix. In the NSBM, the labels for the layers and layer-level groups are sampled from a model consisting of a label-only NDP coupled with independent draws from a beta distribution to form the group-connection probabilities in each layer. This corresponds to drawing $L$ times, once for each layer, from a mixture of GEM distributions with atoms and masses sampled from GEM distributions. This mixture model determines the community $z_{\ell}\in\mathbb{N}$ of layer $\ell$, and assigns to it a class-specific GEM distribution from which the node communities $\xi_{i\ell}$ of each node $i$ in that layer are drawn. A collection of connectivity parameters $\{\eta_{xyz_{\ell}}\}_{x,y\in\mathbb{N}}$ for the layer-community SBM in then associated with each layer, and a connection between nodes $i$ and $j$ in layer $\ell$ is then formed with probability $\eta_{\xi_{i\ell}\xi_{j\ell}z_{\ell}}$. The key distinction between our work and Amini et al. (2024) and Josephs et al. (2023) is the goal of obtaining a global clustering of each node. Furthermore, we develop a nonparametric framework that allows for the incorporation of covariate information. In our model, we allow for node-specific mixtures of the same GEM distributions, with the mixture probabilities $\bm{\tau}_{i}$ driven by nodal covariates. That is, the atoms of each mixture component are the same, but the mixture probabilities differ, being informed by the features. This choice ensures that the clustering of each nodes is comparable and so we obtain a global grouping.
@@ 3 Inference via variational Bayesian methods
The normalising constant for the distribution in (3.1) is not available in closed form, therefore it is necessary to resort to sampling or approximation methods to study the posterior distribution. In particular, we use variational Bayesian (VB) inference to obtain an analytic approximation to a posterior distribution (Wang and Blei, 2019). VB transforms an inference problem into an optimisation problem, which allows for parallelisation and mini-batching (Blei et al., 2017), making it more scalable to large datasets than Markov Chain Monte Carlo (MCMC). In a VB approach, a family of distributions $\mathcal{F}$ is posited on the parameter space of $\Theta=\{\rho,\gamma^{\prime},\varphi,\varphi^{0},\bm{\sigma}^{2},z,\bm{w}\}$, which is partitioned into $S$ disjoint subsets, and one seeks to select the component $q^{\ast}(\Theta)\in\mathcal{F}$ that is closest to the true posterior $p(\Theta\mid X,A)$ in the sense of the Kullback-Leibler (KL) divergence. A mathematically convenient choice for $\mathcal{F}$ is the mean-field variational family:
=$\mathcal{F}=\left\{q:q(\Theta)=\prod_{s=1}^{S}q_{s}(\Theta_{s})\right\},$ (3.2)

### v2

@@ Abstract
This is a pre-print of an article published in the Journal of Complex Networks. The final authenticated version is available online at: https://doi.org/10.1093/comnet/cnaf049.
@@ 1 Introduction
Modern statistics has seen a growing interest in the use of network models to mathematically describe complex relationships between a large number of connected entities. Network data is often highly complex, with multiple types of connections existing between the nodes in the network. To model such heterogeneity of connection types, multiplex networks have been developed as compact representations, with multiple layers being considered simultaneously, each describing different connection types between its nodes. Examples of applications include neuroimaging, where regions of interest in the brain represent nodes and layers correspond to different patients (Paul and Chen, 2018); actions of individuals on a social media platform, where layers might represent “follow” or “retweet” interactions (Greene and Cunningham, 2013); or economic data, where layers capture different types of financial exposure (Poledna et al., 2015). For a survey of multiplex network applications, see, for example, Kivelä et al. (2014).
@@
There has been work to extend single-layer clustering methodology to the multiplex setting, much of which considers variants of a multiplex SBM. Broadly speaking, the approaches can be divided into those that attempt primarily to cluster the layers, and potentially the nodes therein (Stanley et al., 2015; Reyes and Rodriguez, 2016; Paul and Chen, 2018; Jing et al., 2021; Young et al., 2022; Josephs et al., 2025; Zheng et al., 2024), and those that seek only to infer either an across-layer community structure, or within-layer node groups without clustering the layers (Amini et al., 2024; Berlingerio et al., 2011; Kuncheva and Montana, 2015). To the best of our knowledge, a methodology that provides both an across-layer, or global, clustering of each node alongside a layer-level, or local, clustering has not yet been proposed in the literature. In this paper, we aim to address this gap by developing a framework for hierarchical clustering of nodes in a multiplex network that allows heterogeneous clustering across layers, infers latent groups at local and global levels, and incorporates nodal covariate information to guide clustering. This particular type of hierarchical clustering finds relevant applications across several fields. In economics, global clustering can describe trade patterns between entities (Fagiolo et al., 2010). In biology, it helps understand how genes or proteins participate in different biological pathways while also inferring overarching functional groupings (Barabasi and Oltvai, 2004).
As discussed in Miller and Harrison (2018), it is important when modelling to understand whether to treat the number of groups is unknown but finite, or as infinite. Broadly speaking, within Bayesian hierarchical modelling, these settings are treated using parametric and nonparametric models, respectively. However, it should be noted that nonparametric models are also frequently employed to tackle the former setting. In the parametric case, an unknown but finite number of groups is often treated by running an inference procedure multiple times and using criteria such as the Akaike information criterion (AIC; Akaike, 1974) or the Bayesian information criterion (BIC; Schwarz, 1978) to select the optimum number post hoc. Green (1995) developed reversible-jump MCMC (RJMCMC) as an algorithm that selects the number of groups in a single-pass, but it can be very difficult to implement in practice (Miller and Harrison, 2018). Methods have been developed for constructing good reversible jump moves (Brooks et al., 2003; Hastie and Green, 2012), but they add high complexity to the model. Priors can also be placed on the number of components to develop more efficient algorithms Miller and Harrison (2018). To tackle an infinite number of groups, several nonparametric models have been proposed, with two widely used being the hierarchical Dirichlet process (HDP; Teh et al., 2006) and the nested Dirichlet process (NDP; Rodríguez et al., 2008). The infinite nature of these models allows for the selection of the number of latent groups in a single pass. MCMC can be used for these models, broadly using marginalisation Maceachern (1994); Escobar and West (1995), truncation (Ishwaran and James, 2001; Ishwaran and Zarepour, 2002), or RJMCMC. Variational inference is also popular Blei and Jordan (2006), adding underestimation of posterior variance at the gain of a large computational speed-up.
@@
With the exception of Reyes and Rodriguez (2016); Josephs et al. (2025) and Amini et al. (2024), most existing methods assume a known and finite number of communities for layers and nodes. To the best of our knowledge, there are no existing methodologies that have the flexibility of allowing for both an infinite number (or unknown but finite number) of communities and nodal covariate information simultaneously. Within the network literature, approaches for dealing with an unknown or number of nodes usually rely on Bayesian nonparametric priors; for example, Reyes and Rodriguez (2016) modify the infinite relational model of Kemp et al. (2006) to cluster networks with similar community structure. Similarly, the hierarchical stochastic blockmodel (HSBM; Amini et al., 2024) employs an HDP to infer layer-level communities whilst sharing information between the layers. The method proposed in this work is closely related to the nested stochastic blockmodel (NSBM) of Josephs et al. (2025), which uses an NDP to provide a clustering of layers and of the nodes within those layers. The nonparametric prior allows for the number of layer communities and communities within layers to be unspecified and automatically selected. However, the NSBM does not provide a global clustering of nodes, nor is it adapted for the inclusion of nodal covariate information, which we address within the model proposed in this work.
@@ 2 Hierarchical multiplex stochastic blockmodels
We consider a directed multiplex network with $L$ layers and $N$ nodes. We assume that all layers share the same node set $\mathcal{V}:=[N]$, where we write $m]=\{1,\dots,m\},\ m\in\mathbb{N}$. Each layer $\ell\in[L]$ consists of a graph $\mathcal{G}_{\ell}=(\mathcal{V},\mathcal{E}_{\ell})$, where $\mathcal{E}_{\ell}\subseteq\mathcal{V}\times\mathcal{V}$ is a layer-specific edge set, and we say that $(i,j)\in\mathcal{E}_{\ell}$, $i,j\in\mathcal{V}$, if nodes $i$ and $j$ form a connection of type $\ell$. The full multiplex network is denoted by $\mathcal{G}=(\mathcal{V},\{\mathcal{E}_{1},\dots,\mathcal{E}_{L}\})$. Each graph $\mathcal{G}_{\ell}$ could equivalently be represented by its adjacency matrix $A_{\ell}=\{A_{\ell ij}\}_{i,j=1}^{N}\in\{0,1\}^{N\times N}$, where $A_{\ell ij}=\mathbb{I}_{\mathcal{E}_{\ell}}\{(i,j)\}$ and where $\mathbb{I}_{\cdot}\{\cdot\}$ denotes the indicator function. The multiplex network $\mathcal{G}$ can then be represented by an adjacency tensor $A=\{A_{\ell ij},\ \ell\in\{1,\dots,L\},\ i,j\in\{1,\dots,N\}\}\in\{0,1\}^{L\times N\times N}$.
We model the behaviour of each layer by conditionally independent stochastic blockmodels (SBMs; Holland et al., 1983). To parameterise the SBM of layer $\ell\in[L]$, we assume that each node $i\in\mathcal{V}$ belongs to a group $z_{\ell i}\in\mathbb{N}$. Note that, in a Bayesian nonparametric fashion, we consider the number of clusters to be not fixed a priori, and it is instead treated as a random quantity that can grow with the number of nodes and layers. We model the number of groups in the network as an infinite quantity. Given this layer-level group $z_{\ell i},\ \ell\in[L],\ i\in[N]$, the adjacency matrix $A_{\ell}$ is modelled as
@@
and we write $\gamma=\{\gamma_{ks}\}$ and $\gamma^{\prime}=\{\gamma^{\prime}_{ks}\}$ for the matrices storing the parameters. Additionally, we assume that a vector $\bm{x}_{i}\in\mathbb{R}^{P}$ of covariates, which is assumed to be completely observed, is associated to each node $i\in\mathcal{V}$. We use the covariates $\bm{x}_{i}$ to construct a the node-specific probability vectors $\bm{\tau}_{i}$ in (2.3) for each $i\in\mathcal{V}$. To allow for learning of an unbounded number of communities, we utilise a stick-breaking representation (Sethuraman, 1994) for each $\bm{\tau}_{i}$, which ensures that $\sum_{k=1}^{\infty}\tau_{ik}=1$, where $\tau_{ik}$ is the probability that node $i$ belongs to global group $k$. In particular, we employ the probit stick-breaking construction of Chung and Dunson (2009) and Rodriguez and Dunson (2011). In this model, $\bm{\tau}_{i}=(\tau_{i1},\tau_{i2},\dots)^{\intercal}$ is obtained as
@@
$A_{\ell ij}\mid z_{\ell i},z_{\ell j},\rho_{z_{\ell i}z_{\ell j}}$ $\sim\text{Bernoulli}(\rho_{z_{\ell i}z_{\ell j}}),$ $\text{for all }i,j\in\mathcal{V},\;i\neq j,\;\ell\in[L],$ (2.6)
$\rho_{km}$ $\sim\text{Beta}(\alpha_{0},\beta_{0}),$ $\text{for all }k,m\in\mathbb{N},$ (2.7)
$z_{\ell i}\mid w_{i},\bm{\gamma}_{w_{i}}$ $\sim\text{Categorical}(\bm{\gamma}_{w_{i}}),$ $\text{for all }i\in\mathcal{V},\;\ell\in[L],$ (2.8)
$w_{i}\mid\bm{\tau}_{i}$ $\sim\text{Categorical}(\bm{\tau}_{i}),$ $\text{for all }i\in\mathcal{V},$ (2.9)
$\tau_{ik}$ $=\Phi\{\bm{x}_{i}^{\intercal}\bm{\varphi}_{k}\}\prod_{\ell=1}^{k-1}\left(1-\Phi\left\{\bm{x}_{i}^{\intercal}\bm{\varphi}_{\ell}\right\}\right),$ $\text{for all }i\in\mathcal{V},\ k\in\mathbb{N},$ (2.10)
$\bm{\varphi}_{k}\mid\bm{\varphi}^{0}_{k},\sigma_{k}^{2}$ $\sim\text{Normal}(\bm{\varphi}^{0}_{k},\sigma_{k}^{2}I_{P}),$ $\text{for all }k\in\mathbb{N},$ (2.11)
$\bm{\varphi}^{0}_{k}$ $\sim\text{Normal}(\bm{\mu},I_{P}),$ $\text{for all }k\in\mathbb{N},$ (2.12)
$\sigma_{k}^{2}$ $\sim\text{Inverse\text{-}Gamma}(\nu_{0},\omega_{0}),$ $\text{for all }k\in\mathbb{N},$ (2.13)
$\bm{\gamma}_{k}$ $\sim\text{GEM}(\eta_{0}),$ $\text{for all }k\in\mathbb{N}.$ (2.14)
@@
Figure 2: Toy example of the group structure of nodes in a multiplex network with $L=3$ layers corresponding to the illustration in Figure 1(a), with the grey global group as circles and the white as diamonds. There are two global groups, indicated by the shape of the nodes, and three layer groups, indicated by the colours. Probabilities of layer group assignments given the global groups are indicated in the row vectors.
@@
On the other hand, the NSBM is used in Josephs et al. (2025) for clustering of layers in a multiplex network where each layer follows an SBM. Layers are assigned to a community which determines its connection probability matrix. In the NSBM, the labels for the layers and layer-level groups are sampled from a model consisting of a label-only NDP coupled with independent draws from a beta distribution to form the group-connection probabilities in each layer. This corresponds to drawing $L$ times, once for each layer, from a mixture of GEM distributions with atoms and masses sampled from GEM distributions. This mixture model determines the community $z_{\ell}\in\mathbb{N}$ of layer $\ell$, and assigns to it a class-specific GEM distribution from which the node communities $\xi_{i\ell}$ of each node $i$ in that layer are drawn. A collection of connectivity parameters $\{\eta_{xyz_{\ell}}\}_{x,y\in\mathbb{N}}$ for the layer-community SBM is then associated with each layer, and a connection between nodes $i$ and $j$ in layer $\ell$ is then formed with probability $\eta_{\xi_{i\ell}\xi_{j\ell}z_{\ell}}$. The key distinction between our work and Amini et al. (2024) and Josephs et al. (2025) is the goal of obtaining a global clustering of each node. Furthermore, we develop a nonparametric framework that allows for the incorporation of covariate information. In our model, we allow for node-specific mixtures of the same GEM distributions, with the mixture probabilities $\bm{\tau}_{i}$ driven by nodal covariates. That is, the atoms of each mixture component are the same, but the mixture probabilities differ, being informed by the features. This choice ensures that the clustering of each nodes is comparable and so we obtain a global grouping.
@@ 3 Inference via variational Bayesian methods
The normalising constant for the distribution in (3.1) is not available in closed form, therefore it is necessary to resort to sampling or approximation methods to study the posterior distribution. In particular, we use variational Bayesian (VB) inference to obtain an analytic approximation to the posterior distribution (Wang and Blei, 2019). VB transforms an inference problem into an optimisation problem, which allows for parallelisation and mini-batching (Blei et al., 2017), making it more scalable to large datasets than Markov Chain Monte Carlo (MCMC). In a VB approach, a family of distributions $\mathcal{F}$ is posited on the parameter space of $\Theta=\{\rho,\gamma^{\prime},\varphi,\varphi^{0},\bm{\sigma}^{2},z,\bm{w}\}$, which is partitioned into $S$ disjoint subsets, and one seeks to select the component $q^{\ast}(\Theta)\in\mathcal{F}$ that is closest to the true posterior $p(\Theta\mid X,A)$ in the sense of the Kullback-Leibler (KL) divergence. A mathematically convenient choice for $\mathcal{F}$ is the mean-field variational family:
$\mathcal{F}=\left\{q:q(\Theta)=\prod_{s=1}^{S}q_{s}(\Theta_{s})\right\},$ (3.2)
@@
MCMC can be used as an alternative to target the posterior in Dirichlet process models. Broadly, approaches fall into three categories: methods that marginalise out infinite-dimensional parameters (Maceachern, 1994; Escobar and West, 1995); methods that approximate via truncation (Ishwaran and James, 2001; Ishwaran and Zarepour, 2002); and reversible-jump MCMC (RJMCMC) methods (Green and Richardson, 2001; Dahl, 2003; Jain and Neal, 2004). Marginalisation relies on the Pólya urn scheme, which, as discussed in Rodríguez et al. (2008), becomes problematic in the nested setting due to the need to compute complex predictive distributions. While RJMCMC could in principle be applied, constructing efficient trans-dimensional proposals for nested structures is challenging; as shown by Rannala and Yang (2013), split moves that introduce many parameters jointly can result in low acceptance rates and poor mixing. For the HDP, Amini et al. (2019) developed an exact slice sampler, later applied in Amini et al. (2024), but no such analogue is available for the nested case. Truncation-based methods are possible, with Josephs et al. (2025) considering four different Gibbs samplers, but with our addition of covariates and no efficient sampling scheme, computation time quickly becomes a limiting factor. We therefore adopt variational inference in this work.